### v2

# Optimal Bound for PCA with Outliers using Higher-Degree Voronoi Diagrams

Abstract
In this paper, we introduce new algorithms for Principal Component Analysis (PCA) with outliers. Utilizing techniques from computational geometry, specifically higher-degree Voronoi diagrams, we navigate to the optimal subspace for PCA even in the presence of outliers. This approach achieves an optimal solution with a time complexity of $n^{d+\mathcal{O}(1)}\text{poly}(n,d)$. Additionally, we present a randomized algorithm with a complexity of $2^{\mathcal{O}(r(d-r))}\times\text{poly}(n,d)$. This algorithm samples subspaces characterized in terms of a Grassmannian manifold. By employing such sampling method, we ensure a high likelihood of capturing the optimal subspace, with the success probability $(1-\delta)^{T}$. Where $\delta$ represents the probability that a sampled subspace does not contain the optimal solution, and $T$ is the number of subspaces sampled, proportional to $2^{r(d-r)}$. Our use of higher-degree Voronoi diagrams and Grassmannian based sampling offers a clearer conceptual pathway and practical advantages, particularly in handling large datasets or higher-dimensional settings.
1 Introduction
Principal Component Analysis (PCA) is one of the most widely used techniques of dimensionality reduction algorithms in machine learning. In PCA, the goal is to find the best low-rank approximation of a given matrix. This can be conveniently achieved using singular value decomposition (SVD) of the given matrix [12]. SVD decomposes the matrix into three matrices: a left singular matrix, a diagonal matrix of singular values, and a right singular matrix. The low-rank approximation can be obtained by keeping only the top $r$ singular values and their corresponding singular vectors, where $r$ is the desired rank of the low-rank approximation. This is shown in  (1).
$\begin{split}&\texttt{Input:}\>\text{X}\in\mathcal{M}_{n\times d}(\mathbb{R}),r\in\{1,2,\cdots,d\}\\ &\arg\min_{L}||\text{X}-L||_{F}^{2}\ \\ &\texttt{Subject}\>\texttt{to:}\>rank(L)\leq r\end{split}$ (1)
Here $||\text{A}||_{F}^{2}=\sum_{i,j}{a_{ij}^{2}}$ is the square of the Frobenius norm of matrix A.
PCA has many applications in machine learning, such as image processing [16, 4], signal processing [7, 22], neuroscience [8, 19], and quantitative finance [24, 30]. It is often used to reduce the dimensionality of high-dimensional data while preserving the most important information. This leads to improvement in the efficiency and accuracy of machine learning models that work with large high-dimensional data.
In the presence of outliers, the classical PCA algorithm may not work very well, as they can significantly affect the results. The PCA with Outliers problem attempts to improve the results of PCA by removing some data points that are considered outliers. The goal is to recover a low-rank matrix with sparse errors. Here, the assumption is that the data points lie near an $r$-dimensional subspace, and the remaining points are distributed randomly, possibly as a result of distorted observations. This problem can be formulated as follows: given a matrix $\text{X}\in\mathcal{M}_{n\times d}(\mathbb{R})$, where $n$ is the number of data points and $d$ is the dimensionality of the data, and a parameter $r$ representing the rank of the low-rank matrix, the goal is to find a low-rank matrix $L$ and a sparse error matrix $S$ such that $\text{X}=S+L$, where $rank(L)\leq d-r$ and $||S||_{0}\leq k$, where $k$ is a parameter representing the number of outliers to be removed, and $||\cdot||_{0}$ denotes the number of zero entries in a matrix, as shown in (2).
$\begin{split}&\texttt{Input:}\>\text{X}\in\mathcal{M}_{n\times d}(\mathbb{R}),r\in\{1,2,\cdots,d\},\\ &k\in\{1,2,\cdots,n\}\\ &\texttt{Output:}\arg\min_{S,L}||\text{X}-S-L||_{F}^{2}\\ &\texttt{Subject}\>\texttt{to:}\>rank(L)\leq r,||S||_{0}\leq\mathtt{k}\end{split}$ (2)
The PCA with Outliers problem, or robust PCA, has many applications in various fields of research. It can be used to remove noise or artifacts from images in medical imaging [27, 2], detect anomalies or outliers in financial data in portfolio management [10], identify suspicious activities in video data [6], and remove noise or distortions from signals in speech recognition [18]. These applications demonstrate the versatility and usefulness of the PCA with Outliers problem in various fields of research.
Our Results. The contribution of this paper is as follows. First, we propose a new algorithm for PCA with Outliers with time complexity $n^{\mathcal{O}(d)}\text{poly}(n,d)$, which fills the gap identified in [26] as this is the optimal bound. To achieve this, we leverage higher-degree Voronoi diagrams, which are defined in the next section. Second, we develop a randomized algorithm with time complexity $2^{O(r(d-r))}\text{poly}(n,d)$. This algorithm provides an optimal solution for PCA with Outliers with a probability of at least $(1-\delta)^{T}$, where $\delta=\mathbb{P}(\text{Ball}(L,\alpha))$ and $T\propto 2^{r(d-r)}$. Here, $\alpha$ is a constant parameter, and $T$ is the number of subspaces sampled. Also, $\delta$ is defined as the probability of an open ball centered around point $L$ with radius $\alpha$ in the Grassmannian manifold, all to be defined in the later sections. As $\alpha$ decreases, $T$ increases exponentially, ensuring a high success probability. This approach offers a more intuitive solution compared to [26], incorporating an effective sampling technique to accurately hit the optimal subspace. Specifically, we employ a probabilistic method to sample subspaces from the Grassmannian manifold, ensuring that with high probability, the sampled subspaces are close to the optimal solution.
1.1 Related works
In [28], the authors formulated the PCA with Outliers as the optimization problem shown in (3).
$\begin{split}&\min_{L,S}\|L\|_{*}+\lambda\|S\|_{1}\\ &\texttt{subject to:}\quad X=L+S\end{split}$ (3)
Where $\|\cdot\|_{*}$ denotes the nuclear norm of a matrix, which is the sum of its singular values, and $\|\cdot\|_{1}$ denotes the $\ell_{1}$ norm of a matrix, which is the sum of the absolute values of its entries. The parameter $\lambda$ controls the trade-off between the low-rank and sparse components of the data matrix. The PCA with Outliers problem can be solved using a variety of convex optimization techniques, such as the Alternating Direction Method of Multipliers or the Augmented Lagrangian Method.
The paper [29] proposes the Outlier Pursuit method to solve this convex problem. This algorithm is based on the observation that the sparse matrix $S$ can be decomposed into a sum of sparse vectors, each of which corresponds to an outlier. It iteratively identifies and removes the outliers from the data matrix, and then computes the low-rank component of the remaining data using PCA. The outliers are then added back to the low-rank component to obtain the final estimate of the data matrix. The mentioned algorithm requires the sparsity level to be known a priori, which can be a limitation in practice. Additionally, the algorithm assumes that the noise in the data matrix is Gaussian, which may not be true in practice.
To overcome the shortcomings of the previous methods, [9] proposed the Manipulator Pursuit algorithm, which is more robust and flexible. Manipulator Pursuit can handle a wider range of outlier models and noise distributions. It does not require the sparsity level to be known a priori as it uses a data-driven approach to estimate the sparsity level and a nuclear norm regularization term to encourage low-rank solutions.
In [26], the authors analyze the computational complexity of the PCA with Outliers problem. The authors refine the lower bound on the time complexity of the problem, showing that it is $\Omega(n^{d})$. The authors conclude that the lower bound is tight up to a polynomial factor, and also propose an algorithm solving this problem in $n^{\mathcal{O}(d^{2})}{\text{poly}(n,d)}$ using tools from real algebraic geometry. To be explicit, the result is summarized in the following proposition.
Proposition 1.
([5], Theorem 13.22) Consider $V$ as an algebraic set in $\mathbb{R}^{d}$ of real dimension $d^{\prime}$, defined by $Q(X_{1},\ldots,X_{d})=0$, where $Q$ is a polynomial in $\mathbb{R}[X_{1},\ldots,X_{d}]$ of degree at most $b$. Let $\mathcal{P}\subset\mathbb{R}[X_{1},\ldots,X_{d}]$ be a finite set of $s$ polynomials, each $P\in\mathcal{P}$ also having a degree at most $b$. Define $D$ as the ring generated by the coefficients of $Q$ and the polynomials in $\mathcal{P}$. An algorithm exists that, given $Q$, $d^{\prime}$, and $\mathcal{P}$, computes a set of points satisfying every non-empty cell of $V$ over $\mathcal{P}$. This algorithm requires at most $s^{d^{\prime}}b^{O(d)}$ arithmetic operations in $D$.
Proposition 2.
([26], Theorem 1) Solving PCA with Outliers is reducible to solving $n^{\mathcal{O}(d^{2})}$ instances of PCA.
Proof.
Consider an algebraic set $W=\mathbb{R}^{(d-r)\times d}$. For every $V$ in $W$, we define $S_{V}$ consists of $k$ non-zero rows, selected from $X$ based on those with the highest value of $\|x_{i}-\text{proj}_{V}x_{i}\|^{2}_{F}$, equivalent to $\|Vx^{T}_{i}\|^{2}_{F}$. Meanwhile, $L_{V}$ represents the orthogonal projection of the rows of $(X-S_{V})$ onto $V$.
Now, let’s contemplate the set of polynomials $\mathcal{P}=\{P_{i,j}\}_{1\leq i<j\leq n}$ defined over $W$, where
$P_{i,j}(V)=\|Vx^{T}_{i}\|_{F}^{2}-\|Vx^{T}_{j}\|_{F}^{2}.$
The $k$ farthest points and the matrix $S_{V}$ remain consistent across all cells in the partition of $W$ over $\mathcal{P}$. Therefore, it suffices to select a point $V$ from each cell, calculate the outlier matrix $S_{V}$, and solve PCA for $(X-S_{V},r)$.
Combining all steps, their algorithm functions as outlined below:
  1. 1.
    Utilize the procedure outlined in Proposition 1 to acquire a point $V_{C}$ from each cell $C$ of $W$ over $\mathcal{P}$.
  2. 2.
    For each $V_{C}$, compute the optimal $S_{V_{C}}$ by selecting the $k$ rows of $X$ with the highest value of $||V_{C}x^{T}_{i}||_{F}^{2}$. Construct the instance $(X-S_{V_{C}},r)$ for PCA.
  3. 3.
    The solution to the original instance of PCA with Outliers is the optimal solution among all the solutions obtained from the constructed PCA instances.
By running this algorithm, we can reduce the problem of PCA with Outliers to solving $\binom{n}{2}^{(d-r)d}2^{O(d)}$ instances of PCA. ∎
Proposition 3.
[26] There is no $\omega$-approximation algorithm with running time $f(d)\cdot N^{{o}(d)}$ for any $\omega\geq 1$ and computable function $f$ unless Exponential Time Hypothesis (ETH) fails, where $N$ is the bit size of the input matrix $X$.
In [11] new algorithms for the PCA with Outliers problem are proposed. The authors first present a ($1+\epsilon$)-approximation algorithm for the problem, which has a time complexity of $n^{\mathcal{O}(\epsilon^{-2}\cdot r\log r)}$. Then they presented a randomized algorithm for the problem, which outputs the correct answer with probability $1-\delta$ in time $2^{\mathcal{O}(r(\log k+\log\log n)(\log k+\log\log n+\log 1/\delta))}\text{poly}(n,d,1/\delta)$ under the $\alpha$-gap assumption. This assumption generally guarantees that the outliers must have a certain distance to the expected lower-dimensional subspace.
2 Preliminaries
We use $[n]$ to denote the set ${1,2,\cdots,n}$ for a positive integer $n$. The distance of vector $x$ from subspace $V$ denoted by $\mathtt{dist}(x,V)$, is defined as $\inf_{v\in V}||x-v||$. While $\|\cdot\|_{1}$ denotes the $\ell_{1}$ norm of a matrix, which is the sum of the absolute values of its entries. Let $X\subset\mathbb{R}^{d}$ be a set of size $n$, we define $\mu(X)=\frac{1}{n}\sum_{x\in X}x$ as the mean of the vectors in $X$. In this paper, we consider the PCA with Outliers problem, which takes an instance $(X,r,k)$ as input, where $X$ is a subset of $\mathbb{R}^{d}$ of size $n$, and $r,k\in\mathbb{N}$. The goal is to find a $r$-dimensional subspace $L$ that captures the most important information contained in the data points of $X$ while ignoring up to $k$ outliers. These outliers, which are rows of $X$, will be denoted by a (sparse) matrix $S$ of the same size, where only the rows corresponding to the inliers are replaced with zeros. Throughout the paper, we will use matrix form and linear subspace spanned by rows of that matrix, interchangeably.
2.1 Probability Measure on Rank $r$ Subspaces
In this subsection, we establish a probability measure on the set of all rank $r$ subspaces of an $d$-dimensional real vector space. This set is formally represented by the Grassmannian manifold [^1] See [21] for more details on Grassmannian manifolds and their uses. $\text{Gr}(r,d)$, a fundamental object in differential geometry and algebraic topology.
The Grassmannian $\text{Gr}(r,d)$ is the space of all $r$-dimensional subspaces of $\mathbb{R}^{d}$. This manifold is both compact and connected, with its dimension given by:
$\dim\text{Gr}(r,d)=r(d-r).$
The Grassmannian is endowed with a natural topology and a differential structure, making it a Riemannian manifold. This structure allows for the application of various geometric and analytic techniques to study its properties.
To define a measure on $\text{Gr}(r,d)$, we utilize the Haar measure [^2] See [17] for definition and properties of the Haar measure.. The Haar measure is a unique, translation-invariant measure on the manifold, invariant under the action of the orthogonal group $O(d)$. This invariance is crucial as it ensures that the measure does not change under orthogonal transformations, preserving the intrinsic geometry of the manifold. The Haar measure is finite due to the compactness of the Grassmannian manifold, making it suitable for normalization.
To transform the Haar measure into a probability measure, we normalize it such that the total measure of the entire Grassmannian is 1. Let $\mu_{\text{Haar}}$ denote the Haar measure on $\text{Gr}(r,d)$. The normalized probability measure $\mu$ is then defined by:
$\mu(A)=\frac{\mu_{\text{Haar}}(A)}{\mu_{\text{Haar}}(\text{Gr}(r,d))},$
for any measurable subset $A\subset\text{Gr}(r,d)$. This normalization ensures that $\mu$ is a probability measure, i.e., $\mu(\text{Gr}(r,d))=1$.
Definition 1.
Formally, we define the probability measure $\mu$ on the Grassmannian manifold $\text{Gr}(r,d)$ as follows:
$\mu:\mathcal{B}(\text{Gr}(r,d))\to[0,1],$
where $\mathcal{B}(\text{Gr}(r,d))$ denotes the Borel $\sigma$-algebra [^3] For a comprehensive understanding of Borel $\sigma$-algebra, readers may look into the relevant sections in [25]. on $\text{Gr}(r,d)$, and
$\mu(A)=\frac{\mu_{\text{Haar}}(A)}{\mu_{\text{Haar}}(\text{Gr}(r,d))}.$
This construction ensures that $\mu$ is a well-defined probability measure, allowing for a normalized way to measure subsets of the Grassmannian manifold. This enables probabilistic analysis on the set of all rank $r$ subspaces of $\mathbb{R}^{d}$, facilitating the study of random subspaces.
By using this probability measure, we can rigorously analyze the likelihood of certain subspaces containing desired properties, such as minimizing projection distances for PCA. This approach is fundamental in developing algorithms that leverage random sampling of subspaces, providing a theoretical foundation for their effectiveness and efficiency.
Lemma 1.
The volume of the entire Grassmannian, denoted as $\text{Vol}(\text{Gr}(r,d))$, is given by:
$\text{Vol}(\text{Gr}(r,d))=\frac{\prod_{i=1}^{r}\left(\pi^{(d-i)/2}/\Gamma((d-i)/2+1)\right)}{\prod_{i=1}^{r}\left(\pi^{(r-i)/2}/\Gamma((r-i)/2+1)\right)\prod_{i=1}^{d-r}\left(\pi^{(d-r-i)/2}/\Gamma((d-r-i)/2+1)\right)}.$
Proof.
The volume of the Grassmannian $\text{Gr}(r,d)$ is derived from the volumes of the associated manifold and orthogonal groups. The Grassmannian $\text{Gr}(r,d)$ can be seen as the quotient of the orthogonal group $O(d)$ by the product of the orthogonal groups $O(r)$ and $O(d-r)$. Therefore, its volume can be expressed in terms of these groups’ volumes.
In case of the orthogonal group $O(d)$, the volume is given by:
$\text{Vol}(O(d))=\frac{2^{d(d-1)/2}\pi^{d(d+1)/4}}{\prod_{j=1}^{d}\Gamma(j/2)}.$
In case of the manifold $V_{r}(\mathbb{R}^{d})$, the space of orthonormal vectors in $\mathbb{R}^{d}$, the volume is given by:
$\text{Vol}(V_{r}(\mathbb{R}^{d}))=\frac{\pi^{r(d-r)/2}}{\prod_{i=1}^{r}\Gamma((d-i+1)/2)}.$
The Grassmannian can be viewed as the quotient of the manifold by the orthogonal group $O(r)$:
$\text{Gr}(r,d)=\frac{V_{r}(\mathbb{R}^{d})}{O(r)}.$
Hence, the volume of $\text{Gr}(r,d)$ is:
$\text{Vol}(\text{Gr}(r,d))=\frac{\text{Vol}(V_{r}(\mathbb{R}^{d}))}{\text{Vol}(O(r))}.$
Substituting the volumes from the orthogonal group and manifold, we get:
$\text{Vol}(\text{Gr}(r,d))=\frac{\frac{\pi^{r(d-r)/2}}{\prod_{i=1}^{r}\Gamma((d-i+1)/2)}}{\frac{2^{r(r-1)/2}\pi^{r(r+1)/4}}{\prod_{j=1}^{r}\Gamma(j/2)}}.$
Simplifying this expression, we obtain:
$\text{Vol}(\text{Gr}(r,d))=\frac{\pi^{r(d-r)/2}\prod_{j=1}^{r}\Gamma(j/2)}{2^{r(r-1)/2}\pi^{r(r+1)/4}\prod_{i=1}^{r}\Gamma((d-i+1)/2)}.$
Further simplification leads to:
$\text{Vol}(\text{Gr}(r,d))=\frac{\prod_{i=1}^{r}\left(\pi^{(d-i)/2}/\Gamma((d-i)/2+1)\right)}{\prod_{i=1}^{r}\left(\pi^{(r-i)/2}/\Gamma((r-i)/2+1)\right)\prod_{i=1}^{d-r}\left(\pi^{(d-r-i)/2}/\Gamma((d-r-i)/2+1)\right)}.$
Thus, we have derived the volume of the Grassmannian manifold $\text{Gr}(r,d)$. ∎
This measure offers a precise method for conducting probabilistic analyses of subspaces, which is essential for the effectiveness of algorithms that utilize random subspace sampling.
2.2 Voronoi diagrams
Consider a space $X$ (e.g. $\mathbb{R}^{d}$ or a subset of it), a set of objects $S$ (e.g. a subset of $X$), and a distance function $d:X\times S\to\mathbb{R}$ on this space. A Voronoi diagram is a partitioning of $X$ into subsets each closest to the given set of objects $S$. Each element of the set $S$ such as $s_{i}$ has a corresponding region, called a Voronoi cell, consisting of all the elements of the space $X$ closer to that element $s_{i}$ than to any other element $s_{j\neq i}$ in the set. Voronoi diagrams have many applications in various fields [3, 23]. An example of a Voronoi diagram with a data set of size 15 is depicted in Figure 1.
Figure 1: A Voronoi diagram with 15 data points that partition the 2D plane. The query point labeled with $P$ is closer to the point $X_{9}$ than any other points in the dataset, a property that all points surrounding the region of $X_{9}$ share as well.
In the context of the PCA with Outliers problem, we use Voronoi diagrams to assign each data point to a subspace based on its distance to a set of candidate subspaces. This leads to solving the problem more efficiently by reducing the search space and avoiding the need to compute the distance to all possible subspaces. In the following, we present the formal definition of Voronoi diagrams.
Definition 2.
(Voronoi Diagrams) Let $\mathcal{S}\subset\mathbb{R}^{d}$ be a set of $n$ vectors in a $d$-dimensional space. Voronoi diagram is defined as the partitioning $V\subset\mathcal{P}(\mathbb{R}^{d})$ of $\mathbb{R}^{d}$ such that for a vector $s\in S$:
$\forall x,y\in\mathbb{R}^{d}\quad x\sim y\iff\arg\min_{s\in S}\|x-s\|=\arg\min_{s\in S}\|y-s\|$
where $\mathcal{P}(A)$ is the power set of $A$.
Definition 3.
Let $S$ be a set of $n$ subspaces of $\mathcal{L}(\mathbb{R}^{d})$ of rank $k$ and $X$ be the set of all rank $k$ subspaces of $\mathcal{L}(\mathbb{R}^{d})$. If we define distance $Gr_{k}:X\times S\to\mathbb{R}$ as the usual distance function of the Grassmannian, we can define the partitioning $V\subset\mathcal{P}(\mathcal{L}(\mathbb{R}^{d}))$ of $\mathcal{L}(\mathbb{R}^{d})$ such that for a subspaces $Y\in S$:
$\forall U,W\in\mathcal{L}(\mathbb{R}^{d})\quad U\sim W\iff\arg\min_{Y\in S}Gr_{k}(U,Y)=\arg\min_{Y\in S}Gr_{k}(W,Y)$
Although a normal Voronoi cell is defined as the set of elements closest to a single object in S, a $\mathtt{k}$th-degree Voronoi cell is defined as the set of elements having a particular tuple of $k$ objects in $S$ as their nearest neighbors. Higher-degree Voronoi diagrams also subdivide space.
Higher-degree Voronoi diagrams can be generated recursively. To generate the $\mathtt{k}$th-degree Voronoi diagram from set $S$, start with the $\mathtt{(k-1)}$th-degree diagram and replace each cell generated by $T=\{s_{1},s_{2},\cdots,s_{k-1}\}$ with a Voronoi diagram generated on the set $S\setminus T$.
Definition 4.
(Higher Degree Voronoi Diagrams) Let $\mathcal{S}\subset\mathbb{R}^{d}$ be a set of $n$ vectors in a $d$-dimensional space. We define the $\mathtt{k}$th-degree Voronoi diagram as a partitioning induced by a function $V:\mathcal{P}^{k}(\mathcal{S})\to\mathcal{P}(\mathbb{R}^{d})$ such that:
$V(U)=\bigg{\{}\>\>x\>\>\bigg{|}\>\>\max_{u\in U}d(x,u)\leq\min_{t\in\mathcal{S}\setminus U}d(x,t)\>\>\land\>\>d(x,u_{1})\leq d(x,u_{2})\leq\cdots\leq d(x,u_{k})\bigg{\}}$
where $\mathcal{P}^{k}(A)$ is the set of all $\mathtt{k}$-tuples with elements from $A$ and $d(x,y)$ is the Euclidean distance function.
Definition 5.
Let $\mathcal{S}\subset\mathbb{R}^{d}$ be a set of $n$ vectors in a $d$-dimensional space. We define the $\mathtt{k}$th-degree Voronoi diagram of hyperspaces with respect to vectors $\mathcal{S}$ as a partitioning induced by a function $V:\mathcal{P}^{k}(\mathcal{S})\to\mathcal{P}(\mathcal{L}(\mathbb{R}^{d}))$ such that:
$V(U)=\bigg{\{}\>\>W\>\>\bigg{|}\>\>\max_{u\in U}d(W,u)\leq\min_{t\in\mathcal{S}\setminus U}d(W,t)\>\>\land\>\>d(W,u_{1})\leq d(W,u_{2})\leq\cdots\leq d(W,u_{k})\bigg{\}}$
where $\mathcal{P}^{k}(A)$ is the set of all $\mathtt{k}$-tuples with elements from $A$ and $d(x,y)$ is the Euclidean distance function.
An arrangement of hyperplanes in a $d$-dimensional space $\mathbb{R}^{d}$ is a subdivision of the space into distinct regions or cells created by a finite set of hyperplanes. Formally, given a set $H$ of $n$ hyperplanes, the arrangement of $H$ divides $\mathbb{R}^{d}$ into convex polyhedral cells of various dimensions, defined by the intersections and complements of these hyperplanes. Each cell represents a maximal connected subset of $\mathbb{R}^{d}$ that does not intersect any hyperplane from the set. Arrangements play a crucial role in computational geometry and combinatorics as they provide a framework for analyzing spatial subdivisions. Specifically, arrangements can be used to define Voronoi diagrams. In a Voronoi diagram, the space is divided into regions based on the distance to a specific set of points. Each region, or Voronoi cell, consists of all points closer to one generating point than to any other. By interpreting the bisectors between generating points as hyperplanes, the Voronoi diagram can be viewed as a special case of an arrangement where the hyperplanes are equidistant boundaries between pairs of points. This interpretation allows the powerful combinatorial and geometric tools developed for arrangements to be applied to the study and computation of Voronoi diagrams.
Proposition 4.
([13], Theorem 3.3) Let $H$ be a set of $n$ hyperplanes in $\mathbb{R}^{d}$, for $d\geq 2$. Then there is an algorithm that constructs the arrangement of $H$ in $O(n^{d})$ time, and this is optimal.
Theorem 1.
Let $\mathcal{S}\subset\mathbb{R}^{d}$ be a set of $n$ vectors in a $d$-dimensional space. Construction of the $\mathtt{k}$th-degree Voronoi diagram of hyperspaces with respect to vectors $\mathcal{S}$ as defined in definition 5 is computable in time $n^{O(d)}$.
Proof.
We are going to formulate the boundary of each cell of $\mathtt{(n-1)}$th-degree diagram with a set of linear equations which is going to result in $n^{2}$ hyperplanes. using proposition 4, we are able to compute the expected diagram in $n^{O(d)}$. $\mathtt{k}$th-degree Voronoi diagram is simply constructible by omitting the last $n-k$ indices of each cell.
For every two elements $s_{i}$ and $s_{j}$ from $\mathcal{S}$, Let $P_{ij}$ be a function on the set of all subspaces of the vector space $\mathcal{L}(\mathbb{R}^{d})$ to real numbers $\mathbb{R}$ which measures the difference in the distance of $s_{i}$ and $s_{j}$ to a given subspace $W$. So, if the projection operator on $W$ is $P_{W}$ then we can define this as follows:
$P_{ij}(W)=\|P_{W}\cdot s_{i}\|-\|P_{W}\cdot s_{j}\|$
By sorting the values of the $P_{ij}$’s for a specific $W$, we can determine the corresponding cell for this subspace. To elaborate further, since our focus lies only on the sequence of vector distances from $\mathcal{S}$ to $W$, the signs of the equations $P_{i,j}$ suffice. Consequently, a Voronoi cell can be represented by a sign condition based on these equations. Thus, by the continuity of the equations $P_{i,j}$ within the specified domain, we can deduce that a subspace $W$ lies on the boundary of a Voronoi cell only when there exist $i$ and $j$ such that $P_{i,j}(W)=0$. As we have $n^{2}$ equations with a unique rank $d-1$ solution, we have $n^{2}$ hyperspaces and their arrangments construct the expected diagram. This results in a $(n^{2})^{d}=n^{O(d)}$ time algorithm. ∎
Figure 2 shows a Voronoi diagram of the furthest subspace of the 2nd-degree for 8 points within a two-dimensional space, projected onto one-dimensional subspaces.
Figure 2: Voronoi diagram of the furthest subspace of the 2nd-degree for 8 points in the two-dimensional space on one-dimensional subspaces. The projection of subspaces for the Voronoi diagram of the furthest subspace is shown on the unit circle (smaller circle) and the image of subspaces for the Voronoi diagram of the furthest subspace for the second furthest subspace is marked on the circle with radius 2 (larger circle). Space coloring indicates different cells and subspaces of the same color indicate the same cell. For example, in the above figure, 7 cells are specified, considering that 4 of these eight points belong to exactly one subspace, the number of possible cells is far less than all the possible combinations as outlier data.
3 Optimal PCA with Outliers
In this section, we present our algorithm for solving the PCA with Outliers problem in $n^{d+\mathcal{O}(1)}\text{poly}(n,d)$ time. The solution involves obtaining a better subspace embedding for a matrix by removing its outliers. The challenge is that we do not know the actual inlier matrix to compute the scores, as an arbitrary set of rows of the given matrix might be outliers. Nevertheless, we can narrow down the available choices such that no possibility for a given combination of outliers, exists.
In earlier work [26], all possible combinations were generated using methods from real algebraic geometry to model the valid outliers set, and then each cell was tried as a possible candidate. The basic idea behind our method is similar to [26], but we construct a higher-degree Voronoi diagram out of the given dataset in order to model the outliers set, instead of incorporating algebraic methods.
Theorem 2.
The PCA with Outliers problem is reducible to solving $n^{d+\mathcal{O}(1)}$ instances of PCA.
Proof.
Note that by proposition 3, if such an algorithm exists, it should be optimal. In each step, we pick a possible $\mathtt{k}$-tuple of data and label them as outliers, then compute PCA for the rest and measure the relevant error, and finally, depending on the least measured error, we get the output. Nevertheless, the actual problem is the number of possible $\mathtt{k}$-tuples of outliers. In order to specify each of these possible $\mathtt{k}$-tuples, we need to find their corresponding Voronoi cell. However, there is no need to obtain the exact Voronoi diagram as we only need to know if a set of data can be grouped as outliers, which can be determined by the emptiness of the corresponding cell. Algorithm 1 illustrates the above steps.
Algorithm 1 Computing PCA with Outliers using higher-degree Voronoi Diagram
1:procedure PCA($X_{n\times d}$,r,k)
2:     Compute the $\mathtt{(n-k)}$th-degree Voronoi diagram of the columns of $X_{n\times d}$ according to def 5
[…]
4:     $\hat{\textbf{L}}\leftarrow$ corresponding $k$-tuples of cells in L
5:     $\text{solution}\leftarrow\arg\max_{S\in\hat{\textbf{L}}}\text{PCA}(X\setminus S,r,k)$
6:     return solution
7:end procedure
To show the correctness of our algorithm, we need to show that all possible sets of outliers are explored. Notice that a given subset $S$ of size $k$ is valid as outliers if and only if there is a subspace $L$ of $rank(k)$ with $\mathtt{dist}(x,L)<\mathtt{dist}(s,L)$ for all $x\in X\setminus S$ and $s\in S$. Thus a given subset $S$ of size $k$ is valid as outliers if and only if there is a Voronoi cell corresponding to $X\setminus S$, which proves the correctness of our algorithm.
Since the number of generated Voronoi cells is bounded as stated in theorem 1, the total running time of the above algorithm is $n^{d+\mathcal{O}(1)}\text{poly}(n,d)$ which closes the gap in [26].
4 Randomized Algorithm
In this section, we present a randomized algorithm for solving the PCA with Outliers problem. We begin by introducing the concept of the $\alpha$-gap constant, which is a crucial component of our algorithm.
Lemma 2.
($\alpha$-Gap Constant) For a set $\text{X}\subset\mathbb{R}^{d}$, there exists an $\alpha>0$ such that for an optimal solution of PCA with Outliers, involving a reduced matrix $L$ and a set of outliers $S$, we have for all $s\in S$ and $x\in\text{X}\setminus S$,
$\mathtt{dist}(s,L)>(1+\alpha)\mathtt{dist}(x,L).$
Proof.
Let $L$ be the optimal subspace for this problem. Since this is the optimal subspace for the inliers, any point classified as an outlier must have a greater distance to $L$ compared to the inliers. Otherwise, if an outlier was closer to $L$ than some inliers, it would be considered part of the main structure captured by $L$ rather than an outlier. Thus define $d_{1}$ as the distance of the $(n-k)$th farthest vector to $L$ and $d_{2}$ as the distance of the $(n-k+1)$th farthest vector to $L$ (i.e., the farthest inlier and the nearest outlier). We can define $\alpha$ as the ratio between these two distances:
$\alpha=\frac{d_{2}-d_{1}}{d_{1}}.$
Using this constant, we provide a lower bound on the probability that a Voronoi cell contains the optimal subspace. This demonstrates that sampling multiple subspaces uniformly can yield the optimal result with high probability.
Next, we provide background information on Grassmannian manifolds and subspace projections to lay the structure of our algorithm. We prove a lemma concerning the distance between subspaces in a real vector space and how to extract an upper bound on the distance of one point to these subspaces. This allows us to apply a concept similar to the triangle inequality to the set of subspaces and vectors.
Lemma 3.
Let $U$ and $V$ be two subspaces of the same rank in an $d$-dimensional real vector space, and let $x$ be a vector in this space. If the distance between $U$ and $V$ is $\mathtt{dist}(U,V)$(according to the Grassmannian metric), and the distance between $x$ and $U$ is $\mathtt{dist}(x,U)$, then the maximum distance of $x$ from $V$ is $\mathtt{dist}(x,U)(1+\mathtt{dist}(U,V))$.
Proof.
Let $P_{U}(x)$ and $P_{V}(x)$ denote the orthogonal projections of $x$ onto the subspaces $U$ and $V$, respectively. By definition, the distance from $x$ to $U$ is given by:
$\mathtt{dist}(x,U)=\|x-P_{U}(x)\|.$
The Grassmannian distance between the subspaces $U$ and $V$ ($\mathtt{dist}(U,V)$) is related to the largest principal angle $\theta_{\max}$ between them, such that:
$\mathtt{dist}(U,V)=\sin\theta_{\max}.$
To find the maximum distance from $x$ to $V$, consider the point $u=P_{U}(x)$. The distance from $u$ to $V$ is maximized when we project $u$ onto $V$ in the direction corresponding to $\theta_{\max}$. Thus, the distance from $u$ to its projection onto $V$ can be expressed as:
$\|u-P_{V}(u)\|=\|u\|\sin\theta_{\max}=\mathtt{dist}(x,U)\sin\theta_{\max}=\mathtt{dist}(x,U)\mathtt{dist}(U,V).$
By the triangle inequality, the distance from $x$ to $V$ can then be bounded as:
$\|x-P_{V}(x)\|\leq\|x-P_{U}(x)\|+\|P_{U}(x)-P_{V}(u)\|.$
Substituting the known distances, we get:
$\|x-P_{V}(x)\|\leq\mathtt{dist}(x,U)+\mathtt{dist}(x,U)\mathtt{dist}(U,V)=\mathtt{dist}(x,U)(1+\mathtt{dist}(U,V)).$
We now give another lemma regarding the maximal ball around a subspace, which preserves the relative distances of vectors from that subspace in the Grassmannian manifold.
Lemma 4.
Let $x$ and $y$ be two vectors in an $d$-dimensional real vector space, and let $V$ be a rank $r$ subspace of this space such that $(1+\alpha)\cdot\mathtt{dist}(x,V)\leq\mathtt{dist}(y,V)$ for some $\alpha>0$. Then, the maximal open ball in the Grassmannian manifold $\text{Gr}(r,d)$ centered around $V$ such that for every subspace $U$ in this ball, $\mathtt{dist}(x,U)\leq\mathtt{dist}(y,U)$ has radius $\alpha$.
Proof.
Consider the Grassmannian manifold $\text{Gr}(r,d)$, which is the set of all $r$-dimensional subspaces of $\mathbb{R}^{d}$. Let $\mathtt{dist}(x,U)$ denote the Euclidean distance from a vector $x$ to a subspace $U$. Given the subspace $V$ and vectors $x$ and $y$ such that $(1+\alpha)\cdot\mathtt{dist}(x,V)\leq\mathtt{dist}(y,V)$, we aim to find the maximal open ball around $V$ within which for every subspace $U$, the inequality $\mathtt{dist}(x,U)\leq\mathtt{dist}(y,U)$ holds.
The distance between subspaces in the Grassmannian manifold is related to the principal angles between them. For any subspace $U$ close to $V$, the distances $\mathtt{dist}(x,U)$ and $\mathtt{dist}(y,U)$ can be bounded by Lemma 3. Specifically, for a subspace $U$ within a distance $\mathtt{dist}(U,V)$ in the Grassmannian from $V$, the distances are given by:
$\mathtt{dist}(x,U)\leq\mathtt{dist}(x,V)(1+\mathtt{dist}(U,V))$
and
$\mathtt{dist}(y,U)\leq\mathtt{dist}(y,V)(1+\mathtt{dist}(U,V)).$
To ensure $\mathtt{dist}(x,U)\leq\mathtt{dist}(y,U)$, we require:
$\mathtt{dist}(x,V)(1+\mathtt{dist}(U,V))\leq\mathtt{dist}(y,V)(1+\mathtt{dist}(U,V)).$
Given the condition $\mathtt{dist}(y,V)\geq(1+\alpha)\mathtt{dist}(x,V)$, substituting this into the inequality, we get:
$\mathtt{dist}(x,V)(1+\mathtt{dist}(U,V))\leq(1+\alpha)\mathtt{dist}(x,V)(1+\mathtt{dist}(U,V)).$
Dividing through by $\mathtt{dist}(x,V)(1+\mathtt{dist}(U,V))$ (which is positive since $\mathtt{dist}(U,V)\geq 0$ and $\mathtt{dist}(x,V)\geq 0$), we obtain:
$1\leq 1+\alpha.$
Therefore, for the inequality to hold for all subspaces $U$ within this distance, the maximum $d$ must satisfy:
$1+\mathtt{dist}(U,V)\leq 1+\alpha.$
Hence,
$\mathtt{dist}(U,V)\leq\alpha.$
Thus, the maximal open ball in the Grassmannian manifold $\text{Gr}(r,d)$ centered around $V$ such that for every subspace $U$ in this ball, $\mathtt{dist}(x,U)\leq\mathtt{dist}(y,U)$ has radius $\alpha$. ∎
Now we explain how to compute the measure of the aforementioned maximal ball in the Grassmannian manifold $\text{Gr}(r,d)$. Since the Grassmannian manifold is endowed with the Haar measure, the measure of the maximal ball is simply the volume of the ball. Given that the radius of the maximal ball is $\alpha$, and the dimension of the Grassmannian manifold is $r(d-r)$, the measure of the ball can be computed using the formula for the volume of a ball in a Riemannian manifold.
In $\text{Gr}(r,d)$, the volume of a ball of radius $\alpha$ centered at a point $V$ is given by:
$\text{Vol}(\text{Ball}(V,\alpha))=C\times\alpha^{r(d-r)},$
where $C$ is a constant depending on the geometry of the Grassmannian manifold. As computing the exact constant $C$ for the volume of a ball in the Grassmannian manifold $\text{Gr}(r,d)$ is quite complex, we provide the next lemma to give a lower bound on this value.
Lemma 5.
Let $\text{Gr}(r,d)$ be the Grassmannian manifold of $r$-dimensional subspaces of an $d$-dimensional real vector space. Consider a maximal open ball in $\text{Gr}(r,d)$ centered at a subspace $V$ with radius $\alpha$. The volume of such a ball can be expressed in terms of the volume of the entire Grassmannian and a proportionality constant that accounts for the curvature and dimensionality of the space.
Let $\text{Gr}(r,d)$ be the Grassmannian manifold of $r$-dimensional subspaces of a $d$-dimensional real vector space. Consider a maximal open ball in $\text{Gr}(r,d)$ centered at a subspace $V$ with radius $\alpha$. The ratio of the volume of such a ball to the volume of the entire Grassmannian manifold satisfies the following lower bound:
$\frac{\text{Vol}\left(\text{Ball}(V,\alpha)\right)}{\text{Vol}\left(\text{Gr}(r,d)\right)}\geq\alpha^{r(d-r)}\left(\frac{2e}{r(d-r)}\right)^{r(d-r)/2}.$
Proof.
The Grassmannian manifold $\text{Gr}(r,d)$ represents all $r$-dimensional subspaces of $\mathbb{R}^{d}$ and has a well-defined volume due to its compactness. The volume of the entire Grassmannian, denoted as $\text{Vol}(\text{Gr}(r,d))$, is given by:
$\text{Vol}(\text{Gr}(r,d))=\frac{\prod_{i=1}^{r}\left(\pi^{(d-i)/2}/\Gamma((d-i)/2+1)\right)}{\prod_{i=1}^{r}\left(\pi^{(r-i)/2}/\Gamma((r-i)/2+1)\right)\prod_{i=1}^{d-r}\left(\pi^{(d-r-i)/2}/\Gamma((d-r-i)/2+1)\right)}.$
To approximate the volume of a ball of radius $\alpha$ in $\text{Gr}(r,d)$ centered at a subspace $V$, we can consider the tangent space at $V$. For small $\alpha$, the ball in $\text{Gr}(r,d)$ behaves similarly to a Euclidean ball in this tangent space, whose dimension is $r(d-r)$. Thus, the volume of such a ball can be expressed as:
$\text{Vol}(\text{Ball}(V,\alpha))=C\alpha^{r(d-r)},$
where $C$ is a constant that depends on the geometry of the Grassmannian manifold.
To estimate the constant $C$, we consider the relationship between the volumes of the orthogonal groups and the structure of the Grassmannian. A lower bound for $C$ can be derived using the known volumes of these spaces. Specifically, the volume of a Euclidean ball in a $r(d-r)$-dimensional space is given by:
$\text{Vol}(\text{Ball}(V,\alpha))=\frac{\left(2\pi\right)^{r(d-r)/2}}{\Gamma\left(\frac{r(d-r)}{2}+1\right)}\alpha^{r(d-r)}.$
Therefore, the measure of the maximal ball of radius $\alpha$ in the Grassmannian manifold $\text{Gr}(r,d)$ can be bounded below by:
$\text{Vol}(\text{Ball}(V,\alpha))\geq\frac{\left(2\pi\right)^{r(d-r)/2}}{\Gamma\left(\frac{r(d-r)}{2}+1\right)}\alpha^{r(d-r)}.$
To further refine this bound, we apply Stirling’s approximation for the Gamma function:
$\Gamma(z)\approx\sqrt{2\pi}\left(\frac{z}{e}\right)^{z}.$
Using this approximation, we get:
$\Gamma\left(\frac{r(d-r)}{2}+1\right)\approx\sqrt{2\pi}\left(\frac{r(d-r)}{2e}\right)^{r(d-r)/2+1}.$
Substituting this into our volume expression, we obtain:
$\text{Vol}(\text{Ball}(V,\alpha))\geq\frac{\left(2\pi\right)^{r(d-r)/2}}{\sqrt{2\pi}\left(\frac{r(d-r)}{2e}\right)^{r(d-r)/2+1}}\alpha^{r(d-r)}=\frac{2^{r(d-r)/2}\alpha^{r(d-r)}}{\left(\frac{r(d-r)}{2e}\right)^{r(d-r)/2+1}}.$
Simplifying this expression:
$\text{Vol}(\text{Ball}(V,\alpha))\geq\frac{2^{r(d-r)/2}\alpha^{r(d-r)}(2e)^{r(d-r)/2+1}}{(r(d-r))^{r(d-r)/2+1}}.$
Thus, the simplified ratio of the measure of the maximal ball to the volume of the entire Grassmannian is:
$\mathbb{P}\left(\text{Ball}(V,\alpha)\right)\geq\alpha^{r(d-r)}\left(\frac{2e}{r(d-r)}\right)^{r(d-r)/2}.$
This provides a clear and robust lower bound for the measure of a maximal ball in the Grassmannian manifold $\text{Gr}(r,d)$. ∎
Now we give the following theorem which computes the measure of the Voronoi cell of an optimal subspace in the PCA with the Outliers problem, to be used later in our algorithm analysis.
Theorem 3.
Let $X\subset\mathbb{R}^{d}$ be a set of vectors in the $d$-dimensional real vector space, and let $L$ be the optimal subspace of rank $k$ for the PCA with Outliers problem. If $VOD(L)$ is the Voronoi cell of this given subspace according to Definition 5, then selecting a rank $k$ subspace uniformly at random has a positive probability of at least:
$\mathbb{P}\left(VOD(L)\right)\geq\alpha^{r(d-r)}\left(\frac{2e}{r(d-r)}\right)^{r(d-r)/2}.$
Proof.
Using Lemma 5 and Lemma 4, the proof is immediate. As shown by Lemma 4, based on the definition of the Voronoi diagram (Definition 5), we have $\text{Ball}(L,\alpha)\subset VOD(L)$. Thus, $\mathbb{P}\left(VOD(L)\right)\geq\mathbb{P}\left(\text{Ball}(L,\alpha)\right)$, completing the claim. ∎
We conclude this section by presenting a theorem that highlights the effectiveness of our proposed randomized algorithm for solving PCA with Outliers problem. This algorithm leverages the intrinsic geometric properties of the Grassmannian manifold and the concept of the $\alpha$-Gap Constant to iteratively sample subspaces and identify outliers, ensuring a high probability of finding the optimal subspace. By analyzing the Voronoi cells associated with these subspaces, we construct a framework that conforms with the algorithm’s performance guarantees. The theorem shows that the algorithm can efficiently produce near-optimal solutions with high probability, making it a powerful tool for analyzing high-dimensional data.
Theorem 4.
Let $X\subset\mathbb{R}^{d}$ be a set of vectors in a $d$-dimensional real vector space. There exists a randomized algorithm that, with probability at least $(1-\delta)^{T}$, produces an optimal solution for the PCA with Outliers problem. Here, $\delta=\mathbb{P}(\text{Ball}(L,\alpha))$, where $\alpha$ is a constant, and $T$ is the number of subspaces sampled. The algorithm runs in time:
$2^{O(r(d-r))}\text{poly}(n,d),$
where $n$ is the number of vectors, $d$ is the dimensionality of the space, and $r$ is the target subspace dimensionality. The term $\text{poly}(n,d)$ accounts for the operations per sample, including PCA computation.
Proof.
Based on Theorem 3, we know that sampling multiple subspaces uniformly at random has a high probability of hitting the Voronoi cell that contains the optimal subspace. The steps of the algorithm are as follows:
Algorithm 2 Randomized PCA with Outliers
1:$\mathcal{T}\leftarrow$ uniformly sample $T$ subspaces of rank $r$
2:Compute the loss for each subspace in $\mathcal{T}$
3:$\hat{L}\leftarrow$ subspace with the minimal loss
4:$S\leftarrow$ set of $r$ vectors with the largest distances to $\hat{L}$
5:Perform PCA on the dataset $X\setminus S$ after removing the identified outliers
6:return Principal components of $X\setminus S$
Intuitively, by uniformly sampling subspaces, we increase the likelihood of selecting a subspace that is close to the optimal subspace. Each sampled subspace has an independent probability of $1-\delta$ of being within the optimal Voronoi cell. Therefore, by repeating this sampling process $T$ times, the probability that at least one sampled subspace lies within the optimal Voronoi cell is given by the complementary probability of all samples failing, which is $(1-\delta)^{T}$. To achieve a high probability of success (say, at least $1-\epsilon$), we want:
$(1-\delta)^{T}\geq 1-\epsilon.$
Taking the natural logarithm of both sides and using the approximation $\ln(1-x)\approx-x$ for small $x$:
$T(-\delta)\approx\ln(1-\epsilon),$
$T\approx\frac{\ln(1-\epsilon)}{-\delta}.$
Given that $\delta$ is small, $\ln(1-\epsilon)$ approximates $-\epsilon$, so we have $T\approx\frac{\epsilon}{\delta}$ which substituting $\delta=\mathbb{P}(\text{Ball}(L,\alpha))$:
$T\approx\frac{\epsilon}{\alpha^{r(d-r)}\left(\frac{2d}{r(d-r)}\right)^{r(d-r)/2}}.$
The above expression shows that $T$ is inversely proportional to the probability measure of the ball, which itself is a function of $\alpha$:
$T\propto\left(\frac{1}{\alpha}\right)^{r(d-r)}.$
As $\alpha$ decreases, the volume of the ball in the Grassmannian manifold also decreases exponentially. Hence, to maintain a high probability of hitting the optimal subspace, the number of samples $T$ must increase exponentially. We get:
$T\propto 2^{r(d-r)}.$
The computational complexity of this algorithm is $O(T)\times\text{poly}(n,d)$, where $T$ is the number of sampled subspaces, $n$ is the number of vectors, and $d$ is the dimensionality of the space. The polynomial term $\text{poly}(n,d)$ accounts for the operations per sample, such as computing the Principal Component Analysis (PCA) on each sampled subspace. This complexity arises from the repeated sampling and evaluation of subspaces, followed by the PCA computation on the reduced dataset. Given $T\propto 2^{r(d-r)}$, the complexity becomes:
$2^{O(r(d-r))}\text{poly}(n,d),$
highlighting that as the dimensionality reduction factor $\alpha$ decreases, the number of required samples $T$ increases exponentially, which in turn increases the overall computational burden. ∎
In summary, our randomized algorithm leverages the structure of the Grassmannian manifold and the properties of subspace projections to efficiently solve the PCA with Outliers problem with high probability. The approach combines theoretical guarantees with practical feasibility, making it a useful method for handling outliers in high-dimensional data.
5 Conclusion
In this paper, we introduced a new algorithm for the problem of PCA in the presence of outliers, leveraging the concept of higher-degree Voronoi diagrams to effectively detect outliers. Our approach builds on previous work by offering an exact solution with a complexity of $n^{d+\mathcal{O}(1)}\text{poly}(n,d)$, which is more efficient than the solutions in [11] and [26]. Furthermore, our method provides a more intuitive framework to solve the PCA with Outliers problem by clearly delineating the roles of data partitioning and outlier detection. We also presented a randomized version of our algorithm, showing that achieving the optimal subspace can be possible through uniform sampling. While this randomized solution is currently less efficient than the one in [11], we suggest that adopting Gaussian sampling could enhance its accuracy and efficiency, as our main future study direction. Another direction is exploring the dual concept of Voronoi diagrams, specifically Delaunay triangulation, which could provide further insights into the PCA with Outliers problem. The constructive algorithm for Delaunay triangulation presented by [1] offers a promising avenue for investigation. Finally, we would like to link our work to other variants of PCA with outliers, namely online PCA [14, 15]. In particular, we are interested in Bandit PCA [20], where there are feedback loops for having better sequential projections, and where data samples are noisy (have outliers).